OpenStudy (anonymous):

solve −4(x + 1) − 3 = −3(x − 4)

OpenStudy (jiteshmeghwal9):

-4(x+1)-3=-3(x-4) -4x-4-3=-3x+12 -4x-7=-3x+12 -7-12=-3x+4x -19=x

OpenStudy (anonymous):

-4x - 4 - 3+3x -12 = 0 == -x - 19 =0 == x = 19

OpenStudy (jiteshmeghwal9):

wrong @pavaneinstine :(